OpenStudy (anonymous):

what is sin 11pi/2 with working? and without using calcu..

OpenStudy (anonymous):

with working?

OpenStudy (anonymous):

pi = 180, substitute, then evaluate, then reduce

OpenStudy (anonymous):

showing the proper solution

OpenStudy (anonymous):

sin(pi/2) is 1 from unit circle. 11pi/2 should be the same except you must find the quadrant so you know the sign

OpenStudy (anonymous):

sin 11pi/2 = sin pi/2 = 1

OpenStudy (anonymous):

it's -1 cause it's not at pi/2 it's actually at 3pi/2

OpenStudy (anonymous):

Moldy84 has the right answer and explanation.

OpenStudy (anonymous):

sorry - u r right its at 3pi/2

OpenStudy (anonymous):

I don't see anything wrong with my approach

OpenStudy (anonymous):

the trigonometric function value also depends on the quadrant in which the angle lies.

OpenStudy (anonymous):

yea, i just dont know how to get 3pi/2...

OpenStudy (anonymous):

spiko, so 11pi/2 is pi/2 multiplied by 11 right?

OpenStudy (anonymous):

that can put it at 90 degrees or 270 depending on how many times it went halfway around the unit circle

OpenStudy (anonymous):

somtimes it lands at the pi/2 position. sometimes it lands at 3pi/2.

OpenStudy (anonymous):

if it lands at the bottom then it's neg, if top then pos

OpenStudy (anonymous):

Or you can think of it this way: pi/2 = 90 degrees, 1 full circle is 360 degree = 4 pi/2; then 8 pi/2 is 2 full circles; then you add 3 pi/2 to get to 11 pi/2

OpenStudy (anonymous):

spiko, how come you didn't like my approach?

OpenStudy (anonymous):

heromiles i butt hurt

OpenStudy (anonymous):

moldy, im supposed to use this method. example: cos (-31pi/6) = - 30pi/6 - pi/6 = - 5pi - pi/6 = - √3/2 but i dont know how to apply it to that qn i asked

OpenStudy (anonymous):

oh you are using reference angles

OpenStudy (anonymous):

ok then. so 11pi/2 = 10pi/2 = pi/2 = 5pi-pi/2

OpenStudy (anonymous):

sorry + not - on that arithmetic

OpenStudy (anonymous):

sin(5pi) should always be zero then pi/2 but that doesn't really help us cause it's positive. it's been a long time since i've done this. forgive me. i always just think of the unit circle like i was explaining:\

OpenStudy (anonymous):

sin (11pi/2) = 12pi/2 - pi/2 = 6pi - pi/2 = 5pi/2 sin 5pi/2 = -1.506

OpenStudy (anonymous):

i'm pretty rusty at these too - is that the correct answer?

OpenStudy (anonymous):

sorry sin 5pi/2 = 1

OpenStudy (anonymous):

thanks guys...i do know how to get the ans from the unit circle but im supposed to just calculate it using the radians & coordinates of (x,y) & without the use of calcu...jimmy supposed to get 3pi/2 = -1...just really duno the correct way of doin it =/

OpenStudy (anonymous):

Here's the way I did it: sin(11pi/2) = sin(11*180/2) = sin (11*90) = sin(990) = sin (990 - 720) = sin (270) = sin (3pi/2) = -1

OpenStudy (anonymous):

jimmy you got it right i think

OpenStudy (anonymous):

Jimmy didn't get it right...The answer is -1, not 1

OpenStudy (anonymous):

yeah he's right

OpenStudy (anonymous):

yep - good thinking hero - no moldy sin 5pi/2 gives 1 while sin 11pi/2 = -1

OpenStudy (anonymous):

6pi = 0 then we take sin of -pi/2 = -1

OpenStudy (anonymous):

sin(6pi)=0 not sure if we can split it like that. but then you ave the angle -pi/2 left and take the sin of that and you get -1

OpenStudy (anonymous):

heromiles got the right answer

OpenStudy (anonymous):

The way I did it is the easiest way to do it IMO

OpenStudy (anonymous):

heromiles way works too. you get it into two angles that you know from the unit circle and take them seperately. which is how spikos homework i think wanted it done

OpenStudy (anonymous):

sorry meant jimmy

OpenStudy (anonymous):

yes - i just noticed i made a mistake in the last line

OpenStudy (anonymous):

= 6pi - pi/2 = = 5.5 pi and sin 5.5pi = -1

OpenStudy (anonymous):

awesome. thank you so much guys ;)

OpenStudy (anonymous):

5.5pi - 2pi - 2pi = 3pi/2

OpenStudy (anonymous):

The way I did it translates to sin(11pi/2 - 4pi) = sin(11pi/2 - 8pi/2) = sin(3pi/2) = -1

OpenStudy (anonymous):

yep

myininaya (myininaya):

sin(11pi/2)=sin(10pi/2+1pi/2)=sin(5pi+pi/2)=sin(5pi)cos(pi/2)+sin(pi/2)cos(5pi) =0*0+1*(-1)=-1